OpenStudy (haleholmes):

Algebra II Help (Medal) Equation below:

OpenStudy (haleholmes):

2/(4+ square root 6) - 2/(4-square root 6)

OpenStudy (michele_laino):

please, you have to razionalize both your terms. For example, I multiply both numerator and denominator of your first term, by 4-sqrt(6): namely: \[\frac{ 2 }{ 4+\sqrt{6} }=\frac{ 2(4-\sqrt{6}) }{ (4+\sqrt{6})(4-\sqrt{6}) }=\]

OpenStudy (michele_laino):

\[=\frac{ 2(4-\sqrt{6}) }{ 16-6 }=\frac{ 2(4-\sqrt{6}) }{ 10 }=\frac{ 4-\sqrt{6} }{ 5 }\]

OpenStudy (michele_laino):

Now, please do the same with your second term, namely multiply both numerator and denominator, by 4+sqrt(6) @haleholmes

OpenStudy (haleholmes):

\[\frac{ 8+\sqrt{12} }{ 16-6 }\]

OpenStudy (michele_laino):

I got: \[\frac{ 8+2\sqrt{6} }{ 10 }\]

OpenStudy (haleholmes):

I was going to post the more simplified one but my computer was being mean. Alright, now I subtract right?

OpenStudy (michele_laino):

yes!

OpenStudy (haleholmes):

one sec

OpenStudy (michele_laino):

please, keep in mind that your second term is, after simplification: \[\frac{ 4+\sqrt{6} }{ 5 }\]

OpenStudy (haleholmes):

\[4 + \sqrt{6}\]

OpenStudy (michele_laino):

please note that you have to do this calculus: \[\frac{ 4-\sqrt{6} }{ 5 }-\frac{ 4+\sqrt{6} }{ 5 }=...\]

OpenStudy (haleholmes):

I have no idea

OpenStudy (michele_laino):

what is the problem?

OpenStudy (haleholmes):

The calculus. What do I do with that?

OpenStudy (michele_laino):

common least multiple is 5, so we have: \[\frac{ (4-\sqrt{6})-(4+\sqrt{6}) }{ 5 }=\frac{ 4-\sqrt{6}-4+\sqrt{6} }{ 5 }=...\]

OpenStudy (michele_laino):

please, what is the result?

OpenStudy (haleholmes):

It looks like the top just cancels out

OpenStudy (michele_laino):

oops I have made an error...

OpenStudy (michele_laino):

\[\frac{ (4-\sqrt{6})-(4+\sqrt{6} )}{ 5 }=\frac{ 4-\sqrt{6}-4-\sqrt{6} }{ 5 }=...\]

OpenStudy (michele_laino):

please, what is the result?

OpenStudy (haleholmes):

Its being annoying one more sec

OpenStudy (michele_laino):